# Unified Biomolecular Trajectory Generation via Pretrained Variational Bridge

### v1

@@ 4.1 Trajectory Generation
To facilitate comparison, the first 10 ns of one selected MD trajectory was included in the analysis for each test system. As presented in Table 1, PVB achieves performance comparable to the baselines across most metrics, with clear superiority in the validity of generated ensembles. Furthermore, compared to the 10 ns MD trajectories, PVB maintains equivalent physical plausibility and significantly outperforms in distributional similarity for TIC and MSM, underscoring its capability to reliably capture the long-term dynamics of large molecular systems.
@@
We further evaluate the model on the mdCATH dataset, which encompasses a wider range of protein domains. Following Janson et al. (2025), we use a 5,049/40 split for training and validation, with 20 domains selected from the original test set for evaluation. Training data pairs $(x_{t},x_{t+\tau})$ are sampled every 5 ns, with the coarsened time step $\tau=1$ ns. After finetuning, we generate 500-step trajectories for each of the 20 test systems, matching the standard trajectory length in mdCATH. For evaluation, four of the five replica MD trajectories per system are used to construct reference statistics, while the fifth serves as the MD oracle for direct comparison with PVB.
@@ B.2 Proposition 2
We first provide the derivation of Equation 10 from Equation 9. Given ${\mathbf{Y}}_{0}=y_{0}$, we consider the following SDEs:
@@
Let ${\mathbb{P}}_{|0}^{*}$ and ${\mathbb{P}}_{|0}^{u}$ the probability measure of Equation 15 and Equation 16, respectively. Assume that the two SDEs admit strong solutions on $[0,1]$. Since the drift terms are both adapted to filtration of ${\mathbf{B}}_{t}$, we can derive the Radon-Nikodym derivative by applying the Girsanov theorem:
$\log\frac{\mathop{}\!\mathrm{d}{\mathbb{P}}_{|0}}{\mathop{}\!\mathrm{d}{\mathbb{P}}_{|0}^{u}}({\mathbf{Y}}_{0:1})=-\int_{0}^{1}u(t,{\mathbf{Y}}_{0},{\mathbf{Y}}_{t})\mathop{}\!\mathrm{d}{\mathbf{B}}_{t}-\frac{1}{2}\int_{0}^{1}\|u(t,{\mathbf{Y}}_{0},{\mathbf{Y}}_{t})\|^{2}\mathop{}\!\mathrm{d}t.$ (17)
@@
$=\mathbb{E}_{{\mathbf{Y}}_{0:1}\sim{\mathbb{P}}^{u}}[\int_{0}^{1}u(t,y_{0},{\mathbf{Y}}_{t})\mathop{}\!\mathrm{d}{\mathbf{B}}_{t}+\frac{1}{2}\int_{0}^{1}\|u(t,y_{0},{\mathbf{Y}}_{t})\|^{2}\mathop{}\!\mathrm{d}t]$ (20)
$=\mathbb{E}_{{\mathbf{Y}}_{0:1}\sim{\mathbb{P}}^{u}}[\frac{1}{2}\int_{0}^{1}\|u(t,y_{0},{\mathbf{Y}}_{t})\|^{2}\mathop{}\!\mathrm{d}t].$ (21)
@@ F.1 Featurization of Proteins
To featurize trajectories of proteins in ATLAS, we select as collective variables the sine and cosine of backbone $\psi$ and $\phi$ torsions and all side-chain torsions, and the Euclidean distances of contacting heavy-atom pairs, which are defined as pairs of C, N, or S atoms separated by more than two bonds in the molecular topology graph and lying within 4.5 Å in the initial conformation of the MD reference trajectory (Yuan et al., 2012). In particular, since each trajectory in mdCATH contains relatively few frames, we instead use only the backbone $\phi$ and $\psi$ torsions as TICA features for this dataset to avoid ill-conditioning. The extracted features are then projected onto the slowest two time-lagged independent components (TIC0 and TIC1) (Pérez-Hernández et al., 2013) using deeptime (Hoffmann et al., 2021), with a lag time of 100 ps for ATLAS and 1 ns for mdCATH.
@@ G.1 Fast-Folding Proteins
[…]

### v2

# arXiv:2602.07588v2[cs.LG] 27 Feb 2026

@@ 4.1 Trajectory Generation
To facilitate comparison, the first 10 ns of one selected MD trajectory was included in the analysis for each test system. As presented in Table 1, PVB achieves performance comparable to the baselines across most metrics, with clear superiority in the validity of generated ensembles. Moreover, compared to 10 ns MD, PVB maintains equivalent physical realism and achieves superior distributional agreement in TIC and MSM feature spaces. This is further evidenced by a higher decorrelation ratio, confirming its ability to resolve slow dynamical modes inaccessible to short MD simulations.
@@
We further evaluate the model on the mdCATH dataset, which encompasses a wider range of protein domains. Following Janson et al. (2025), we use a 5,049/40 split for training and validation, with 20 domains selected from the original test set for evaluation. Training data pairs $(x_{t},x_{t+\tau})$ are sampled every 5 ns, with the coarsened time step $\tau=1$ ns. After finetuning, we generate 500-step trajectories for each of the 20 test systems, matching the standard trajectory length in mdCATH. For evaluation, four of the five replicate MD trajectories per system are used to construct reference statistics, while the fifth serves as the MD oracle for direct comparison with PVB.
@@ B.2 Proposition 2
We first provide the derivation of Equation 10 from Equation 9. Given the same initial state $y_{0}$, we consider the following two SDEs, assumed to admit unique strong solutions on $[0,1]$:
@@
Let ${\mathbb{P}}_{|0}^{*}$ and ${\mathbb{P}}_{|0}^{u}$ be the probability measures of the solutions to Equation 15 and Equation 16, respectively. Since the drift terms are both adapted to filtration of ${\mathbf{B}}_{t}$, we can derive the Radon-Nikodym derivative by applying the Girsanov theorem:
$\log\frac{\mathop{}\!\mathrm{d}{\mathbb{P}}_{|0}^{*}}{\mathop{}\!\mathrm{d}{\mathbb{P}}_{|0}^{u}}({\mathbf{Y}}_{0:1})=-\int_{0}^{1}u(t,{\mathbf{Y}}_{0},{\mathbf{Y}}_{t})\mathop{}\!\mathrm{d}{\mathbf{B}}_{t}-\frac{1}{2}\int_{0}^{1}\|u(t,{\mathbf{Y}}_{0},{\mathbf{Y}}_{t})\|^{2}\mathop{}\!\mathrm{d}t,~{\mathbf{Y}}_{0:1}\sim{\mathbb{P}}_{|0}^{u}.$ (17)
@@
$=\mathbb{E}_{{\mathbf{Y}}_{0:1}\sim{\mathbb{P}}_{|0}^{u}}[\int_{0}^{1}u(t,y_{0},{\mathbf{Y}}_{t})\mathop{}\!\mathrm{d}{\mathbf{B}}_{t}+\frac{1}{2}\int_{0}^{1}\|u(t,y_{0},{\mathbf{Y}}_{t})\|^{2}\mathop{}\!\mathrm{d}t]$ (20)
$=\mathbb{E}_{{\mathbf{Y}}_{0:1}\sim{\mathbb{P}}_{|0}^{u}}[\frac{1}{2}\int_{0}^{1}\|u(t,y_{0},{\mathbf{Y}}_{t})\|^{2}\mathop{}\!\mathrm{d}t].$ (21)
@@ F.1 Featurization of Proteins
To featurize trajectories of proteins in ATLAS, we select as collective variables the sine and cosine of backbone $\phi$ and $\psi$ torsions and all side-chain torsions, and the Euclidean distances of contacting heavy-atom pairs, which are defined as pairs of C, N, or S atoms separated by more than two bonds in the molecular topology graph and lying within 4.5 Å in the initial conformation of the MD reference trajectory (Yuan et al., 2012). In particular, since each trajectory in mdCATH contains relatively few frames, we instead use only the backbone $\phi$ and $\psi$ torsions as TICA features for this dataset to avoid ill-conditioning. The extracted features are then projected onto the slowest two time-lagged independent components (TIC0 and TIC1) (Pérez-Hernández et al., 2013) using deeptime (Hoffmann et al., 2021), with a lag time of 100 ps for ATLAS and 1 ns for mdCATH.
@@ G.1 Fast-Folding Proteins
To better characterize folding dynamics, we employ global structural descriptors as TICA collective variables: the sine and cosine of backbone ($\phi,\psi$) and side-chain torsions, along with $\alpha-C$$\alpha$ distances for residue pairs separated by at least two residues. The lag time for TICA is taken as 20 ns.
@@ Appendix H Computing Infrastructure
Appendix I Disclosure of Large Language Model Usage
Large language models (LLMs) were used solely as general-purpose writing assistants in the preparation of this manuscript. Specifically, LLMs were employed to polish wording, improve clarity, and correct grammar. They were not used for research ideation, methodological design, experimental planning, data analysis, or any scientific decision-making. All technical content, results, and conclusions were developed and verified by the authors, who take full responsibility for the manuscript.